OpenStudy (jjuden):

What is the distance between points P(-2, 1, 3) and Q(-1, 4, -2) to the nearest tenth?

Nnesha (nnesha):

do you know the distance formula ??

OpenStudy (jjuden):

no

OpenStudy (bibby):

do you know the muffin man?

OpenStudy (jjuden):

maybe

OpenStudy (bibby):

\(\large \sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\)

Nnesha (nnesha):

\(\rm\color{green}{o^\wedge\_^\wedge o}\) ;)

OpenStudy (jjuden):

its 5.9 right

OpenStudy (bibby):

hey, those points have 3 coordinates

OpenStudy (jjuden):

P(-2, 1, 3) and point Q(-1, 4, -2) x = 1, y = 3, z = 5 d = (1^2 + 3^2 + 5^2)^-2 d = (35)^-2 d = 5.9

OpenStudy (bibby):

anyhow, wolfram agrees, sqrt(39) or 5.9